# Optimal and Fair Encouragement Policy Evaluation and Learning

### v4

@@ 8 Case Studies
We first include an empirical case study that analyzes heterogeneous treatment effects and compliance effects in Homonoff and Somerville (2021), which demonstrates the importance of disaggregated evaluations of efficacy and takeup. Next, we consider approaches based on constrained policy optimization.
8.1 Case study: Text message reminders
We consider a case study in the context of text message reminders of recertification interviews for SNAP, a benefit program providing electronic transfers for food purchases, formerly known as food stamps. We first discuss background on the data and the context of the text message reminders, which are expected to increase compliance with the benefit recertification process. Next, we describe our focal analysis task: reoptimizing reminders to optimize impacts (encouragement) on attendance of a recertification interview (treatment) and therefore on next-year benefits (outcome). We first provide a descriptive analysis outlining the heterogeneous effects of encouragement on treatment and of treatment on outcomes, with a focus on comparing group differences. Next, we turn to off-policy evaluation and optimization, considering budget allocations and where potential equity concerns may arise in the distribution of final outcomes ($\mathbb{E}[Y(\pi)\mid A=a]$) vs. the distribution of improvements.
8.1.1 Background
Applications for public services are complex. Federal guidelines require an interview, during which a caseworker typically goes over the application and explains to the applicant complex eligibility guidelines and rules, such as idiosyncratic income reporting and household determination guidelines. The interview can be completed in person or over the phone, after which individuals must provide requested documents within ten days. Recent digital innovations (such as GetCalFresh) have improved the application process. Exact operational delivery is the responsibility of state and local administrators, with some variation in procedural features such as flexible callback systems, scheduling of interviews, or central management of call-in interviews by a call center. We focus on the recertification interview: Every year, to renew benefits, households must conduct another interview to recertify their eligibility, in case of changes to their income or household composition. In San Francisco, California, the case study location, interviews were scheduled completely at random over the course of the next month and could be conducted in person or by phone. In-person interviews impose a hassle cost on the applicant, for example, due to transportation or childcare costs, inflexible work schedules, or other precarities experienced by low-income households applying for benefits.
We reanalyze the data from Homonoff and Somerville (2021), focusing on the text message reminders. Homonoff and Somerville (2021) are primarily interested in the effect of the time between the notice of the scheduled interview and the interview itself on individuals’ recertification and downstream outcomes, such as yearly SNAP benefits. Overall, they find positive impacts of the time lapse between the scheduling of the interview and the end of the month when the recertification process should be completed. The data also record the results from a contemporaneous pilot study of text message reminders about the recertification deadline, which is our focus. Individuals could opt into a text message reminder that their recertification was due at the end of the month and directing them to call the San Francisco Human Services Agency (SFHSA) in case of questions. In addition to the interview, recertification requires the individuals gather certain documents to prove eligibility, with intermediate deadlines for these other processes. The reminder about the overall deadline is theorized to improve awareness of time-sensitive aspects of recertification and improve interview compliance. The original paper (Homonoff and Somerville, 2021) finds a positive average treatment effect. Important questions remain, however, about potential heterogeneity in the impacts of the text message reminders on interview completion (and of interview completion on yearlong outcomes). We focus on the question of reoptimizing textmessage nudges. We leave the question of joint optimization of text message reminders and interview time for future work and assume that current scheduling operations remain fixed. The interview date within the next month is assigned completely at random to balance workload. Assuming that individuals arrive randomly throughout the month, it is plausible that the distribution of the time lapse prior to interview date itself remains random.
Text message reminders are a low-touch, tactical lever to reduce behavioral frictions and costs that could lead to missed interviews. If a beneficiary fails to complete the interview within the given 30-day time window, her benefits are terminated, and although households can later reschedule, benefits are not prorated, leading to benefit losses. In our analysis, we focus on potential optimization of text message reminders and equity concerns. Given that text message reminders are unlikely to reduce interview attendance rates, without resource constraints, the results of Homonoff and Somerville (2021) already suggest they should be made available to all. On the other hand, resource constraints are often binding for nonprofits and government, and thus cost optimization can be critical: Eligibility determinations affect wide swathes of the population, and more intensive outreach options could further strain budgets. For example, the Benefits Data Trust (BDT), a longtime service provider, recently declared the end of its operational model, a move speculated to be due in part to high operational costs.
Our methods illustrate the tradeoffs characterizing different potential targeting rules. For example, our methodology allows more precise estimates of the average encouragement effect, which can be used to evaluate the value of the so-called default option question in nudge designs: What are the gains from a move from opting in to opting out of text message reminders? Could the global budget constraints necessary in practice inadvertently introduce inequality?
8.1.2 Model and descriptive statistics: Heterogeneous takeup and efficacy
Next, we describe our specific instantiations of our general encouragement model. We let $R\in\{0,1\}$ denote receipt of a text message reminder, where $R=1$ denotes that a beneficiary received a reminder. We consider the treatment $T\in\{0,1\}$, where $T=1$ indicates attendance of any recertification interview (enrollees can reschedule another interview after missing their first, although this delays benefit receipt). The downstream outcome $Y\in\mathbb{R}^{+}$ that we consider is the cumulative total of benefits received over the next 12 months. Missing an interview delays benefit receipt, but Homonoff and Somerville (2021) find that outcomes are skewed: Some would-be recipients do not further pursue recertification after missing an interview, potentially losing out on thousands of dollars’ worth of benefits. Homonoff and Somerville (2021) establish via linkage to administrative wage data that many of these individuals are indeed SNAP-eligible and thus they rule out explanations based on purely economic models of ordeal targeting, which interpret a beneficiary’s missing an interview to mean that her household does not need benefits. This finding indicates that behavioral frictions in service delivery can have negative effects. In other words, reduction of frictions can improve service delivery.
Descriptive statistics provide evidence on the existence of treatment effect and compliance heterogeneity and potential drivers of inequity in unconstrained optimal encouragement regimes by disaggregating the treatment effect and compliance heterogeneity (vs. the heterogeneity in the treatment effect of the nudge alone). In Table˜3, we display the results of a logistic regression of recommendation on covariates. In the original data, significant explanatory factors include household size, whether the interview was conducted by phone or in person, gender (female), race (nonwhite), age, whether the beneficiary has children, whether the beneficiary receives the maximum benefit amount, and time lapse to the interview date. Next, in Table˜4, we include the regression table from a recommendation-interacted logistic regression of treatment $T$ on $(X,R,R\times X),$ i.e., covariates and interacted interview reminder. We find significant heterogeneity based on age (interview reminders are less efficacious as age increases), while reminder efficacy increases for those without income in the previous quarter (which indicates more need for benefits/a lower opportunity cost of interview attendance) and with years since first SNAP enrollment (which indicates greater familiarity with the recertification process).
8.1.3 Comparison of potential improvements upon the status quo
$p_{1\mid 1}(x,a)-p_{1\mid 0}(x,a)$ $\tau(x)$ $P(T=1\mid X,A)$ $E[Y\mid X,A]$
Correlation -0.08 -0.28 -0.27 0.04
Correlation A=0 0.23 -0.27 -0.21 0.21
Correlation A=1 -0.15 -0.28 -0.27 0.04
Is self-selection sufficient for targeting?
An important question is whether individuals self-select into the reminder on the basis of their expected value of benefits or the particular effect of the reminder on compliance. We leverage the estimated effect heterogeneity and predictive models to unpack this relationship. We consider the Spearman correlation between the predicted probability of enrollment in the reminder, $P(R=1\mid X,A)$, and $p_{1\mid 1}(x,a)-p_{1\mid 0}(x,a)$ (the heterogeneous effect of compliance); $\tau(X)$ (the heterogeneous effect of treatment), $P(T=1\mid X,A)$ (the marginal treatment-enrollment probabilities); and $E[Y\mid X,A]$ (the marginal expected benefits). The first two reflect the effect heterogeneity: A strong positive rank correlation of selection into encouragement (the reminder) with the compliance or heterogeneous treatment effects (or the product thereof) would indicate that self-selection is efficient for recommendation/treatment efficacy. Since this may be difficult to justify a priori, as enrolling individuals do not know this estimated heterogeneity in the effects, we also assess rank correlations of $P(R=1\mid X,A)$ with the marginal treatment and outcomes. We find that, overall, self-selection is weakly negatively correlated with heterogeneous compliance, treatment effects, and marginal compliance. However, these relationships are heterogeneous by group membership: We find a weakly positive correlation for the white subgroup, where self-selection is weakly correlated with heterogeneous compliance effects and marginal outcome levels. However, generally no such relationship exists for nonwhite beneficiaries. This indicates that current self-selection patterns are overall not efficient for the targeting of reminders to impact outcomes and that, in fact, self-selection is usually negatively correlated with the heterogeneity in the treatment effect: Better-resourced households with lower treatment effects (eligible for lower benefits amounts) are somewhat more likely to enroll in recertification reminders. At a high level, this suggests that there is room for improvement upon self-selection into deadline reminders, whether by changing the default option to a reminder opt-out or by means of covariate-conditional targeting under a budget constraint, as we explore next. A recent Federal Communications Commission (FCC) ruling (Federal Communications Commission, 2021) clarifies that “state governments can communicate information to beneficiaries via text messaging without first obtaining their ‘opt-in’ and without facing liability under the TCPA [Telephone Consumer Protection Act of 1991]”.
Comparison of automatic enrollment in reminders (recommend to all) and optimized budgeted allocation
Next, we consider potential improvements from automatic enrollment in reminders (recommend to all) or, under potential budget constraints, budget-optimal targeted allocations. In Figure˜2, we compare all of these potential policies. The y-axis measures different values of $E[Y(\pi)]$ in differently colored series of lines (red, purple, blue): red for the group $A=1$ (nonwhite), $E[Y(\pi)\mid A=1]$; purple for the entire population, $E[Y(\pi)]$; and blue for $A=0$ (white), $E[Y(\pi)\mid A=0]$. We optimize the policies on a 70% training split. The off-policy values are obtained via doubly robust estimation, with logistic regression as the propensity score and gradient-boosted regression on log-transformed outcomes for outcome regression. However, because of the zero-inflated/heavy-tailed outcomes, we find some instability in the off-policy estimates computed on the test set alone. Thus, we pool the data and obtain the plotted estimates of $E[Y(\pi)\mid A=1]$ from the entire dataset.
Next, within each set of colored lines, we compare the following: The solid line ranging over budgets indicates $\mathbb{E}[Y(\pi_{\%})]$, where $\pi_{\%}$ is the optimal policy under a budget corresponding to delivery of recommendations to some percentage of the population. The x-axis shows this percentage ranging from 0% (no one is treated) to 100% (everyone is treated). Next, the set of dotted/dashed lines presents the outcomes under unpersonalized or untargeted policies: The densely dotted line indicates treat-all, $E[Y(1)\mid A=a]$; the dashed line indicates the observed self-selection (and hence averages over outcomes in the dataset within group status); and the sparsely dotted line indicates $E[Y(1)\mid A=a]$. Within a set of colored lines, on the left side, the budget is large, so the value of the targeted allocation, $E[Y(\pi_{\%})]$ (solid line), tracks the value of therecommend-to-all policy, $E[Y(1)]$; while on the right side, the value of small budgets tracks the value of the recommend-to-none policy, $E[Y(0)]$.
First, we note that, for the most part, any additional reminders generally improve outcomes—whether via self-selection, targeted allocation, or recommendation-to-all. However, as we saw in the previous analysis of self-selection, the value under currently observed self-selection does not improve significantly because of inefficient targeting, although there is heterogeneity: Self-selection somewhat improves outcomes for the white group. Of course, this depends on many factors, some of which may shift in different settings and implementations. Race may be correlated with income and digital literacy (enrollment with phone number), and the reminders differed by device. For most budgets covering treatment of at least $25\%$ of the population, we find that targeted allocation improves group-average outcomes significantly over those under current self-selection or all-control (no reminders). To conclude, regarding equity in outcomes, i.e., group-level outcomes $E[Y(1)\mid A=1],E[Y(1)\mid A=0]$, targeted allocation improves upon a self-selection or no-recommendation status quo.
Budget A=0 Improvement A=0 Equiv. Budget A=1
50.0% \$12.22 71.7%
25.0% \$21.97 50.2%
10.0% \$39.37 29.2%
With these findings in mind, we turn to a final investigation: Suppose that we could consider allocations with different budgets per group rather than a global budget. Given that different compliance responses lead to unequal improvements, how much more outreach toward nonwhite populations would be needed under the allocation to equalize the amount of improvement over the no-allocation baseline? We analyze this question in Figure˜4. We now examine policies that consider different thresholds for different groups, thereby resulting in different budgets, $\pi_{\%,A}=\mathbb{I}[(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\tau(X,A)>q_{A}].$ Typically, $q_{A}$ is a quantile of the score distribution and therefore corresponds to some budget percentile. We let $F_{A}(p)=P(p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\tau(X,A)\leq p)$ denote the cumulative distribution function (CDF) of the heterogeneous encouragement score, and therefore, the groupwise budget at some threshold $q_{a}$ is $F^{-1}_{A}(q_{A})$. We now range over these thresholds separately for the different groups on the $x$-axis; the $y$-axis plots the average improvement in the allocated subgroup, $E[p_{1\mid 1}(X,A)-p_{1\mid 0}(X,A))\tau(X,A)\mid A,\pi_{\%,A}=1]$. In Table˜2, we illustrate for a few potential groupwise budget percentages of $A=0$ what improvement is achieved for $A=0$ and what percentage of the budget would be required for $A=1$ to achieve equal improvement. The equivalent budget allocations required to target the same percentage of the group population can differ by 20 to 25 percentage points. At thresholds $q=F_{0}^{-1}(p)$ corresponding to percentiles $p=50\%,75\%$ and $90\%$ of the $A=0$ group, and therefore budgets of $1-p=50\%,25\%$ and $10\%$, we observe average improvements of $\mathdollar 12,\mathdollar 21.97$ and $\mathdollar 39.37$ in the $A=0$ group, and it would require an equivalent budget of $72\%,50\%$ and $29.2\%$ of the $A=1$ population to achieve equal improvements, respectively. Given that 80% of individuals are nonwhite (i.e., the $A=1$ group is four times the size of the $A=0$ group), substantially different groupwise budgets (in terms of number of people covered by the budget allocation) would be required to equalize improvements.
Summary and Conclusion
In this case study, we focus on assessing the heterogeneous compliance and treatment effects of text message reminders about recertification deadlines on 1) completion of any interview and 2) yearly benefits. First, we focus on insights on the self-selection process and heterogeneity in the encouragement and treatment effects that help us understand potential drivers of inequity in targeted allocations. Although we find that the heterogeneity in the effects of the encouragement (text message reminders about the recertification deadline) on treatment (successful completion of an interview) is positively correlated with the heterogeneity in the treatment effects on downstream benefits collected, the strength of this correlation varies by race: It is stronger for white than for nonwhite beneficiaries. This finding is relevant since the naive heterogeneous encouragement effect can be related to the product of heterogeneous encouragement and treatment effects and so the relative strength of this correlation indicates that an approach based purely on ranking the heterogeneous encouragement effects under the efficient budget allocation would result in some white beneficiaries ranking more highly than nonwhite ones.
Figure˜1 and Figure˜3 illustrate that this inequity arises because of the differential effectiveness of the reminder in improving the probability of interview attendance, suggesting either that a design with further communication could be helpful or that other, structural compliance barriers could persist. Proceeding with the analysis of reminder efficacy, we find that while racial disparities in current patterns of self-selection into encouragement are small, self-selection is still inefficient for targeting: Those who would benefit the most from receiving reminders do not sign up for them. We assess budget-optimal targeted allocations that can improve upon the outcome under self-selection with as little as 10% of the budget. The heterogeneous effects of the reminder on interview completion, i.e., the differences that the nudge makes, are smaller for nonwhite beneficiaries, but this group generally receives higher benefits overall. Overall, any efficiency-targeted allocation rule improves groupwise outcomes (i.e., average 12-month benefits for white or nonwhite beneficiaries) relative to those in the no-reminder baseline.
In conclusion, we find evidence of effect heterogeneity that implies that targeted allocations or opt-out reminders can improve significantly upon the outcomes under current self-selection. Text message reminders and communication infrastructure open up a tactical space for operational service design. Investigating potential targeted allocations, we find that they improve equity in final outcome levels but that inequity persists in how much groups’ outcomes improve on average, driven almost entirely by racial gaps in compliance. Although our tools enable us to dig deeper into the efficacy of self-selection and potential alternative allocations and estimate disparities in outcomes, more social science research is necessary to elucidate the root causes. [^1] Commentaries on nudges in other domains (e.g., criminal justice Kohler-Hausmann, 2020) highlight varying explanations consistent with nonresponse. This is especially important because whether the observed disparities in compliance reflect behavioral inattention, trust in government, or structural barriers to compliance with administrative burdens matters for policy implications.
@@ 8.1.3 Comparison of potential improvements upon the status quo
We conduct a case study on a dataset of judicial decisions on supervised release based on risk-score-informed recommendations for supervised release under an electronic-monitoring program (Office of the Chief Judge, 2019a). The Public Safety Assessment Decision-Making Framework (PSA-DMF) uses a prediction of failure to appear for a future court date to inform pretrial decisions, including the decision that we focus on here: supervised release with electronic monitoring, where judges make the final decision (psa, 2016). While the literature on pretrial risk assessment is large, to our knowledge, it is unclear what empirical evidence justifies the release recommendation matrices that have been used in practice to recommend supervised release via electronic monitoring. [^2] We focus on supervised release with electronic monitoring, though the broad term supervised release encompasses substantially different programs nationwide, including access to supportive services and caseworkers, which has been touted as a factor enabling bail reform and release more broadly (Akinnibi and Holder, 2023). There are current policy concerns about disparities in the increasing use of supervised release, given mixed evidence on its outcomes (Office of the Chief Judge, 2019a; Gross, ); e.g., Safety and Justice Challenge (2022) concludes that "targeted efforts to reduce racial disparities [in supervising release] are necessary". We focus on a publicly available dataset from Cook County, Illinois, with information about defendant characteristics, algorithmic recommendations for electronic monitoring, detention/release/supervised release decisions, and failure to appear and other outcomes (Office of the Chief Judge, 2019b). The data were initially used to assess bail reform (Office of the Chief Judge, 2019a). [^3] While in this initial case study we work with publicly available data (Office of the Chief Judge, 2019b), in future work, we will seek more granular data with additional robustness checks to support our substantive conclusions. We offer a more detailed discussion in the appendix, but to summarize, unconfoundedness is likely violated in this case (but can be addressed with standard sensitivity analysis), and some line-level data are aggregated for privacy.
@@
We let $Z\in\{0,1\}$ denote release $(Z=1)$ (with or without conditions). All of our analysis occurs in the $(XZ,AZ,RZ,YZ)$ group, i.e., among the released population only. For brevity, we drop $Z$ in describing the data below. We let $X$ denote covariates (age, top charge category, PSA failure to appear/new criminal arrest (FTA/NCA) score bucket and flag). The (binarized) protected attribute is $A$: race (nonwhite/white) or gender (female/male). The algorithmic recommendation is $R$, a recommendation from the PSA-DMF matrix for supervised release (at any intensity of supervision conditions). The treatment $T$ is whether the individual is released under supervision (at any intensity of supervision conditions). The outcome variable, $Y$, is failure to appear ($Y=1$).
In Figure˜6, we provide descriptive information illustrating heterogeneity (including by protected attribute) in adherence and effectiveness. We observe wide variation in judges’ assignment of supervised release beyond the recommendation. We use logistic regression to estimate outcome and treatment response models. The first figure shows our estimates of the causal effect by gender (with similar heterogeneity by race). The outcome is failure to appear, so negative scores are beneficial. The second figure illustrates the difference in responsiveness: how much more likely decision-makers are to assign treatment when there is than when there is not an algorithmic recommendation to do so. The last figure plots a logistic regression of the lift in responsiveness on the causal effect $\tau(x,a)=\mu_{1}(x,a)-\mu_{0}(x,a)$. We observe disparities in how responsive decision-makers are conditional on the same treatment effect efficacy. This is, importantly, not a claim of animus because decision-makers did not have access to causal effect estimates. Nonetheless, disparities persist.
Figure 6: Policy value $V(\pi^{\lambda})$, treatment value $\mathbb{E}[T(\pi^{\lambda})\mid A=a],$ for AA = race, gender.
@@
In Figure˜6, we highlight results from constrained policy optimization. The first two plots in each set illustrate the objective function value and $A=a$ average treatment cost for $A=$ race (nonwhite/white) and gender (female/male), respectively. We use costs of $100$ for $Y=1$ (failure to appear, 0 for $Y=0$, and $20$ when $T=1$ (set arbitrarily)) and minimize costs. On the x-axis, we plot the penalty $\lambda$ that we use to assess the solutions of Proposition˜10. The vertical dashed line indicates the solution achieving $\epsilon=0$, i.e., parity in treatment takeup. Near-optimal policies that reduce treatment disparity can be of interest given advocacy concerns about how the expansion of supervised release could increase the surveillance of already surveillance-burdened marginalized populations. We see that, indeed, for race, surveillance parity–constrained policies can substantially reduce disparities for nonwhite defendants while not increasing surveillance on white defendants that much: The red line decreases significantly with a low increase of the blue line (and low increases to the objective value). On the other hand, for gender, the opportunity for improvement in the surveillance disparity is much smaller.
Acknowledgments
We thank the San Francisco Human Services Agency for access to the data and Peri Weisberg for data access and discussions.
Appendix A Additional discussion
A.1 Additional discussion on motivating fairness constraints
[Christensen et al., 2020] posits a “human capital catch-22" that certain axes of precarity such as scarcity and health both increase likelihood of requiring access to state assistance while reducing the cognitive resources required to navigate administrative burdens in service delivery. Some strategies noted for reducing administrative burden in public benefit and service programs OMB [2022] include reducing information/learning costs, which can be modeled with encouragement designs and targeted recommendations; or reducing redemption costs, which can be modeled in a setting of scarce resources with constraints and potentially personalized using the methods we develop here.
A.2 Additional related work

### v5

@@ 6 Case Studies
6.1 Why and where the two-stage constrained optimization helps
In Section 5.2, we developed a general two-stage procedure for constrained formulations or Pareto improvement upon prior policies. We highlight the benefits of our two-stage approach in a simulated example, where it operates as effective variance regularization and improves out-of-sample fairness constraint feasibility and Lagrangian regret, especially in finite samples.
Figure 4: Comparison of the two-stage method of Section 6.2 to ablations, as a function of sample size $n$. In the disparity panel, the horizontal line is the nominal fairness target $\epsilon=0.3$, while the dashed curves labeled 1-stage, 2-stage, and two-stage-75 denote the corresponding slack-adjusted effective constraint levels: for the one-stage baseline, $\epsilon_{\mathrm{eff}}^{(1)}=\bigl(\epsilon-z_{0.9}\sigma_{\max}/\sqrt{n}\bigr)_{+},$ and for two-stage methods, obtained from first-stage variance estimates at the effective sample size. The last panel measures Lagrangian regret relative to each method’s own effective constraint levels.
In Figure 4, we compare several ways of solving the same constrained optimization problem. The naive method solves the empirical constrained problem directly, with no extra slack. The one-stage baseline adds a single worst-case variance slack, while two-stage implements our procedure: it uses a first split to identify a candidate solution and nearly binding constraints, then re-solves a localized problem on a second split with variance-based slacks. The two-stage-75 variant uses the same localization idea with larger effective subsamples (75% of the dataset in a split), and Pareto-imp-multiobj denotes the local Pareto-improvement variant around a first-stage optimal policy rather than the global constrained problem. Implementation details are deferred to Appendix Section 13.2. Figure 4 reports, from left-to-right, key metrics measured on an out-of-sample test set: value regret, treatment disparity, coverage of the target fairness constraint, and a variant of Lagrangian regret that accounts for different conservativities due to different effective sample complexities of the one- vs. two-stage methods. The main pattern is that the naive approach is least conservative but often violates the fairness constraint out of sample, whereas the regularized baselines improve feasibility.
Across the simulation settings, the naive approach is least conservative but frequently violates the fairness constraint out of sample, whereas the one-stage and two-stage procedures both improve constraint satisfaction. For the value regret, in general, fairness is imposed at some value loss; there is a trade-off in general since compliance factors into the encouragement score as well. The advantage of the two-stage method is most visible in the moderate-sample regime, where it attains similar or better policy value while delivering tighter control of disparity. At larger sample sizes, this advantage narrows as the cost of sample splitting becomes less worthwhile.
@@ 6.1 Why and where the two-stage constrained optimization helps
We conduct a case study on a dataset of judicial decisions on supervised release based on risk-score-informed recommendations for supervised release under an electronic-monitoring program (Office of the Chief Judge 2019a). The Public Safety Assessment Decision-Making Framework (PSA-DMF) uses a prediction of failure to appear for a future court date to inform pretrial decisions, including supervised release with electronic monitoring. An existing decision-making matrix recommends supervised release based on thresholds of failure-to-appear (FTA) and new-criminal-activity (NCA) predictive risk scores (psa 2016). To our best understanding, such thresholds were not previously optimized on data. [^2] We focus on supervised release with electronic monitoring, though the broad term supervised release encompasses substantially different programs nationwide, including access to supportive services and caseworkers, which has been touted as a factor enabling bail reform and release more broadly (Akinnibi and Holder 2023). These thresholds on predictive risk scores need not align with optimal thresholds on encouragement scores. There are current policy concerns about disparities in the increasing use of supervised release, given mixed evidence on its outcomes (Office of the Chief Judge 2019a, Gross ); e.g., Safety and Justice Challenge (2022) concludes that ”targeted efforts to reduce racial disparities [in supervising release] are necessary”. We focus on a publicly available dataset from Cook County, Illinois, with information about defendant characteristics, algorithmic recommendations for electronic monitoring, detention/release/supervised release decisions, failure to appear and other outcomes (Office of the Chief Judge 2019b).
@@
All of our analysis occurs in the released population only. We let $X$ denote covariates (age, top charge category, PSA failure to appear/new criminal arrest (FTA/NCA) score bucket and flag). The (binarized) protected attribute is $A$: race (nonwhite/white) or gender (female/male). The algorithmic recommendation is $R$, a recommendation from the PSA-DMF matrix for supervised release (at any intensity of supervision conditions). The treatment $T$ is whether the individual is released under supervision (at any intensity of supervision conditions). The outcome variable, $Y$, is failure to appear ($Y=1$).
In descriptive diagnostics included in Section 13.4, we compare the heterogeneity in treatment effects and responsiveness. We find estimated heterogeneity in treatment efficacy and compliance, as well as group-level differences in distributions. We also observe disparities in how responsive decision-makers are to recommendation, conditional on the same treatment effect efficacy. This is importantly not a claim of animus because decision-makers didn’t have access to causal effect estimates. Nonetheless, disparities persist.
Figure 5: Pretrial supervised release case study: policy optimization via enumeration over $\lambda$ in penalty formulation. Local scalarized score $100\,\mathbb{E}[Y(\pi^{\lambda})]+20\,\mathbb{E}[T(\pi^{\lambda})]$ and treatment rate $\mathbb{E}[T(\pi^{\lambda})\mid A=a]$, for AA = race, gender.
@@
In Figure 5, we highlight results from constrained policy optimization. The first two plots in each set illustrate the local scalarized score $100\,\mathbb{E}[Y(\pi)]+20\,\mathbb{E}[T(\pi)]$ and the group-$A=a$ treatment rate for $A=$ race (nonwhite/white) and gender (female/male), respectively. This scalarization is specific to the supervised-release application. On the x-axis, we plot the penalty $\lambda$ that we use to assess the solutions of Proposition 5.2. The vertical dashed line indicates the solution achieving $\epsilon=0$, i.e., parity in treatment takeup. Near-optimal policies that reduce treatment disparity can be of interest given advocacy concerns about how the expansion of supervised release could increase the surveillance of already surveillance-burdened marginalized populations. We see that, indeed, for race, surveillance parity–constrained policies can substantially reduce disparities for nonwhite defendants while not increasing surveillance on white defendants that much: The red line decreases significantly with a low increase of the blue line (and low increases to the local scalarized score). On the other hand, for gender, the opportunity for improvement in the surveillance disparity is much smaller.
Fairness-constrained optimization over interpretable scorecards.
A key challenge in this setting is the lack of overlap in recommendations for supervised release. Prior work (Ben-Michael et al. 2021) addresses this by performing robust optimization over encouragements, effectively extrapolating outcome models. In contrast, our decomposition of encouragement into treatment take-up and treatment efficacy allows us to rely on treatment overlap while restricting robustness to the treatment response to recommendations. This is a milder extrapolation problem, as it can be guided by behavioral structure such as monotonicity.
We optimize over local regions of potential improvement upon the existing status quo policy. We use Proposition 5.5 to optimize over uncertainty sets based on Lipschitz parametric extrapolation of high-overlap regions to low-overlap regions. The status quo policy is given by a decision-making framework on combinations of FTA and NCA scores, and we model $R=1$ if the highest levels of supervision are recommended. For privacy, the predictive risk scores were discretized into score buckets 1-2, 3-4, 5-6 etc. We optimize over the discrete space of interpretable scorecards, i.e. comprising of recommendation decisions for each (FTA, NCA) pair because these are nearly the same form as the current decision-making matrix. We also consider covariate-richer age-split scorecards that expand policy expressivity while remaining interpretable. (Age has widely been empirically shown to be a major driver of discretion in pretrial decision-making).
Figure 6:
To optimize over Lipschitz recommendation response models $p_{1\mid r}(x),$ we use a characterization of Khan et al. (2024) which reduces the Lipschitz uncertainty set on $p_{1\mid r}(x)$ to a per-datapoint interval uncertainty set, compatible with Proposition 5.5. We calibrate the Lipschitz constant $L=0.17$ based on the largest effective Lipschitz constant observed on the good-overlap region. We also impose the additional compliance monotonicity constraint that $p_{1\mid 1}(x)-p_{1\mid 0}(x)\geq 0,\;\forall x$. To preserve interpretability, we optimize over monotone scorecards, i.e. recommendation within one $(FTA,NCA)$ pair requires recommendation for larger $(FTA,NCA)$ pairs in a partial order. This partial order is pre-calculated and imposed as another linear constraint. Finally, under the status quo policy, about $18\%$ are recommended stricter supervised release: we assume this treatment budget holds and impose it as an additional constraint. These structural constraints can be collected via linear polytope constraints on Proposition 5.5. We consider a scalarized outcome with FTA:treatment at a 100:20 ratio. See the appendix Section 13.4.1 for more details and the full formulation.
Figure 6 displays the results, where each “scorecard” has FTA scores on the y-axis, and NCA scores on the x-axis. Even the pooled cost-optimal scorecard substantially reduces worst-case disparity relative to the status quo (0.0246 versus 0.0572). Allowing age-stratified scorecards improves worst-case disparity slightly further, to 0.0229 or 0.0217. Thus, local changes to simple decision matrices can reduce disparity while remaining interpretable. It is important that in this situation, even cost-minimizing scorecards without explicit fairness penalty can improve worst-case disparities relative to the current status quo, subject to our robust Lipschitz extrapolation partial identification intervals. Although our local improvements over interpretable scorecards are more restricted compared to our full policy optimization earlier, we can therefore provide concrete recommendations for local changes to status-quo decision-making matrices that could improve disparities at little cost in FTA. Optimizing over covariate-rich data-driven recommendation policies can mitigate fairness-accuracy concerns by providing Pareto-improving local improvements upon the status quo.
Figure 7: Interpretable supervised release scorecards optimized over uncertainty sets on $\tau$ alone with $\Gamma=1.247$. Worst-case disparities assessed with $L=0.204$.
Remark 6.1 (Robustness under violations of unconfoundedness)
One concern could be potential violations of Section 3 (unconfoundedness), in that decisions about treatment relied on additional information that is unmeasured in our dataset. In Section 13.4.2, we leverage the robust linear program Proposition 5.5 to instead optimize over bounds on heterogeneous treatment effects under bounded violations of treatment unconfoundedness. Further, we audit worst-case disparities of the confounding-robust-optimal policies over the same Lipschitz uncertainty set on overlap. In Figure 7 we find that our main findings hold directionally, in that similar confounding-robust scorecards can robustly control disparities estimated at 0.0263~0.0263, though the bounds result in more conservative assessment of who stands to benefit.
7 Conclusion
Across these case studies, the main managerial implication is that organizations should explicitly model both treatment take-up and treatment efficacy, and audit disparities in each separately, because gaps in either mechanism suggest different managerial responses. In the SNAP setting, our analysis can quantify the benefits and limits of improvements from targeting. Low-touch interview reminders improve outcomes on average, but for some more than others — operational redesign of outreach, access, defaults, or service navigation can ensure that those who stand to gain the most do. Our extension to algorithmic advice in the supervised-release case study illustrates that searching over covariate-rich personalized rules can uncover local Pareto improvements over simpler, non-data-driven status-quo rules, reducing disparity with little increase in objective costs.
Appendix contents.
  • Section 8: Additional Discussion (additional related work; alternative identification)
  • Section 9.1: A simple model with low fairness-value tradeoff
  • Section 10: Additional discussion on method (extrapolation; overlap)
  • Section 11: Additional discussion on general optimization method (constrained optimization; saddle-point algorithm)
  • Section 12: Proofs (identification and characterization; generalization; robust characterization; general algorithm)
  • Section 13: Additional case studies and experimental details (SNAP; simulation; OHP; PSA-DMF)
8 Additional Discussion
8.1 Additional Related work